Solve the system of linear equations by graphing and state the solution. Graph and label at least three points on each line and the intersection point, if it exists. If not state so and why.
4𝑥 − 𝑦 = 7
2𝑥 − 3𝑦 = −9

Answer by mananth(16946)


4𝑥 − 𝑦 = 7
x=1,y=-3 (1,-3)
x=3,y=5(3,5)
x=2,y=1 (2,1)

2𝑥 − 3𝑦 = −9
x=-3,y=1 (-3,1)
x=-6, y=-1 (-6-1)
x=0, y=3 (0,3)

solution (3,5)
.